OpenStudy (ajprincess):

Pls help:) \[\frac{dy}{dx}=xy+x+y+1\] Change this D.E into homogeneous form.

OpenStudy (mimi_x3):

\[\frac{dy}{dx} = xy + y + x + 1\] \[\frac{dy}{dx} = y( x + 1) + (x+1)\] \[\frac{dy}{dx} = (x+1)(y+1)\] \[\frac{dy}{dx} * \frac{1}{(y+1)} = (x+1)\] then integrate both sides?

OpenStudy (ajprincess):

ya by this way we can do it. and I already did it. what I want is to convert that equation in the form M(x,y)dx+N(x,y)dy=0. Is it possible @mimi_x3

OpenStudy (mimi_x3):

exact questions?

OpenStudy (mimi_x3):

equations**

OpenStudy (ajprincess):

homogeneous eqns in the form dy/dx=f(y/x)

OpenStudy (ajprincess):

in some books homogeneous eqns form is given as M(x,y)dx+N(x,y)dy=0 that's y first gav that form

OpenStudy (mimi_x3):

you want to use \[ u = \frac{y}{x}\] then convert it into exact equations? \(M(x,y)dx + N(x,y)dy=0\) <= exact equation

OpenStudy (mimi_x3):

I'm confused..

OpenStudy (yttrium):

What math is this? >:D

OpenStudy (mimi_x3):

differential equations.

OpenStudy (yttrium):

Oh no, I'm not taking up this yet. HAHA. Teach me! :D

OpenStudy (anonymous):